### v2

# Pressure Drop in Non-Spherical Packed Beds: Influence of Geometry and Reynolds Number

Abstract
Understanding fluid flow through porous media with complex geometries is essential for improving the design and operation of packed-bed reactors. Most existing studies focus on spherical packings, having as a consequence that accurate models for irregular interstitial geometries are scarce. In this study, we numerically investigated the flow through a set of packed-bed geometries consisting of square bars stacked on top of each other and arranged in disk-shaped modules. Rotation of each module allows the generation of a variety of geometrical configurations at Reynolds numbers of up to 200 (based on the bar size). Simulations were carried out using the open-source solver OpenFOAM. Selected cases (e.g., $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\mathrm{Re}_{\mathrm{p}}=100,200$) were compared against Particle Image Velocimetry measurements. Results reveal that, based on the relative rotation angle, the realized geometries can be classified as channel-like ($\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and lattice-like, fundamentally altering the friction factor. Furthermore, the maximum friction factor obtained in the creeping regime occurred at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas in the inertial regime, this occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Varying the rotation angle also affects the transition from the viscous to the inertial regime.
keywords:
Friction factor , Non-spherical porous medium , Packed bed , Onset of inertial regime , Non-conformal coupling
journal: Particuology
[…]
[label1]organization=Ruhr University Bochum, addressline=Department of Mechanical Engineering, city=Bochum, postcode=44801, state=North Rhine-Westphalia, country=Germany
1 Introduction
A packed bed consists of a solid packing, typically a particle assembly confined within an enclosure, and passed by a fluid. Packed beds are widely used in petroleum, process, and chemical industry (seckendorff21). Typical examples are catalyst reactors (MiguelGarcia20), where residual methane undergoes total oxidation to produce carbon dioxide and water; catalytic converters for emission control, e.g. (Schnitzlein87), and thermal energy storage systems (RyanAnderson15), or as filters in environmental engineering (ChenWang19). The solid particles are often modelled as monodisperse spheres; in reality, however, they can span wide ranges of sizes and shapes, e.g., cylinders, polyhedra, or arbitrary non-polyhedral forms such as in (Afandizadeh2001; Schlipf2015).
The presence of the solid phase forces the fluid, in general, to adopt complex meandering paths as it permeates the packing, enhancing the heat and mass transfer and facilitating reactions (Dixon20). The gas flow in the interstitial space, which is the focus of this study, is influenced by both micro- and macroscopic parameters of the packing, e.g., particle size and shape (or pore size), bed porosity and permeability, and the operating conditions of the device characterized for example, by the Reynolds number (Kumar23; Abdulmohsin17). Therefore, a clear understanding of the impact of these factors on the flow features is crucial for designing efficient reactors (Elzubeir24).
Due to the complexity of these systems, detailed investigation, both numerical and experimental, covering all relevant scales is not viable. Therefore, numerical models of varying fidelity are generally used in the design process. For example, the whole device can be represented with a fast, one-dimensional model neglecting the geometrical complexity (niedermeier2018; Nash2017, e.g.,). When the problem has to be modelled in two or three dimensions, the flow field can be described at a macroscale level (i.e., using averaged flow properties) using porous medium models (collazo2012; battiato2019; Sadowski2023). This typically involves modelling the drag induced by the porous medium using closures (Whitaker1986; whitaker1996) such as Darcy–Forchheimer (DF). Complementary to these continuum approaches, particle movement and macroscopic parameters of the particle assembly can be tracked by the Discrete Element Method (Cundall1979; illanamahiques2023; Ma2022).
Higher accuracy of representation of the flow field is possible resolving the geometry of the packing, and ensuring a sufficiently accurate spatio-temporal discretisation to capture all relevant scales (Dixon20). Such approach—called Particle-Resolved Direct Numerical Simulation (PRDNS)—offers the most precise representation of both micro- and macroscale flow properties within packed beds. PRDNS can provide detailed data on interphase momentum, energy, and mass transfer allowing to develop, calibrate, and validate closure models for unresolved simulations (Tenneti2014; Sadowski2023b; Sadowski2023). However, such simulation requires high computational effort and ,hence, is typically limited to simple and laboratory-scale configurations (neeraj2023; sadowski24; sadowski2025).
A further current limitation of investigations of particle assemblies and packed beds is the simplification of the particle geometry, whereby particles are generally considered as spherical (e.g., zhu16; dentz22; storm24), even in approximations of industrial devices (neeraj2023; sadowski24), for example in a body-centered cubic (BCC) arrangement. Spherical particles offer practical advantages as the calculation of wall-distances can be enormously simplified and the definition of particle contact regions is relatively straightforward. This is more challenging for irregular shapes (Dixon20; jurtz2019) so that spherical particles allow for efficient contact detection and simulation of full-scale systems (illanamahiques2023a).
However, the higher complexity of the flow field patterns induced by heterogeneous, non-spherical packings can result in large errors in velocity predictions by models describing the flows in terms of averaged velocity (Moghaddam19). Moreover, roding2017 observed that the effective diffusivity in random packings of cuboids, spheres and ellipsoids is significantly influenced by the particle shape at identical porosity. This is troubling from the perspective of unresolved flow modelling, as most models express the drag or diffusivity in terms of macroscopic parameters and would be unable to capture such differences. Standard correlations for pressure drop, such as the Ergun equation (ergun1951) or Kozeny–Carman model (carman1937), were originally developed only for spherical particles. Although these and similar drag laws have, to some extent, been successfully generalized to other geometries (e.g., liu94; duplessis1988; duplessis2008; woudberg2020), to the best of our knowledge, no studies have been conducted to quantify how the variation in packing geometry influences the induced drag at constant porosity for non-spherical packing.
Therefore, this study explores how the variation of the interstitial geometry affects the induced drag in a dense packed bed consisting of square bars. To quantify these effects, we carried out numerical simulations in a set of different geometries with the same porosity ($\phi=0.322$) at Reynolds numbers (based on bulk velocity and bar size) spanning the range between $0.1$ and $200$. Special attention was given to the influence of the geometry on the transition from viscosity- to inertia-dominated flow regimes. Moreover, to ensure the quality of the comprehensive database resulting from the 266 simulations, we carried out a detailed validation of the numerical setup against the Particle Image Velocimetry (PIV) results of christin2025. Furthermore, a thorough mesh sensitivity analysis was performed to assess the viability of the numerical approach.
The remaining part of this work is organized as follows. In Section 2, the geometry of the packed bed and the associated flow conditions are described in detail. In Sections 3 and 4 the mathematical and numerical models used in the study, respectively, are described. Section 5 presents the results with particular emphasis on the non-dimensional parameters in porous media. Finally, we summarize our conclusion in Section 6.
Figure 1: (a) The laboratory-scale packed bed of christin2025. (b) The cross-section of the packed-bed geometry formed from six modules, each rotated by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ relative to the preceding one. (c) Schematic of the geometry of each module showing the circle enclosing the dodecagon defining the outer geometry of the slits. The streamwise direction is oriented perpendicular to the surface of the paper.
Figure 2: Visualizations of the simulation domains (i.e., volumes occupied by the fluid) for each of the studied rotation angles, ranging from $5^{\circ}$ to $90^{\circ}$. The streamwise direction is oriented vertically, with the flow moving from bottom to top.
2 Geometry and Flow Conditions
The geometry investigated in this work is based on a laboratory-scale modular packed-bed reactor designed by christin2025. The reactor, shown in fig. 1, is a column consisting of identical modules and an outlet section which is not considered in this study. Each module is a disk with a thickness of $B=$10\text{\,}\mathrm{m}\mathrm{m}$\$, with $5\text{\,}\mathrm{m}\mathrm{m}$ wide slits cut through, forming prismatic bars with a square cross-section ($B\times B$) within each module. The corresponding dimensions are shown in fig. 1. The geometry of the void space in each module is defined so that the side walls of the slits are lying on a regular dodecagon inscribed in a circle with a diameter $D=$62.12\text{\,}\mathrm{m}\mathrm{m}$\$, which is concentric with the modules outer geometry. The area of this circle defines the bed’s cross-sectional area, leading to a theoretical porosity of $\phi=0.322$.
The disks are stacked vertically and rotated relative to one another, creating a complex void-space geometry between adjacent layers. An example of the resulting geometry, with a rotation angle of $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ between layers, is shown in fig. 1.
The fluid moves through the column in the axial direction. The flow is characterized by a Reynolds number based on the bar width $B$ and an intrinsic average (see eq. 3) of the flow velocity through the packed bed $\langle w\rangle$:
$\mathrm{Re}_{\mathrm{p}}=\frac{\langle w\rangle B}{\nu},$ (1)
where, $\nu$ is the fluid’s kinematic viscosity.
By rotating each module by a fixed angle, different interstitial configurations and flow characteristics in the bed can be investigated. In the present study, $19$ rotation angles and, thereby, 19 geometries, are considered: $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$,$5\text{\,}\mathrm{\SIUnitSymbolDegree}$,\dotsc,$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. In order to accurately represent the macroscopic flow properties of each of packing, it is vital that a periodic geometry is considered (guibert2016; scandelli2022), otherwise a blockage effects would influence the computed values of pressure drop. For this reason, for each rotation angle, such a number of layers is chosen, which results in a periodically repeating geometry. The number of layers corresponding to each rotation angle is provided in table 1.
Figure 2 shows the simulation domains corresponding to each angle. For a small rotation angle (up to $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), the arrangement resembles a channel-like structure, where aligned void spaces create curved flow passages. As $\alpha$ increases above $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, the void spaces become interconnected and the geometry becomes more complex, with the exception of $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. At $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the geometry again exhibits a channel-like character.
3 Mathematical Model
An incompressible flow of a Newtonian fluid with constant and homogeneous density (i.e., neglecting variations due to mixing, reaction, or thermal effects) inside the considered packed bed can be described at the pore level by the Navier–Stokes and continuity equations:
$\frac{\partial u_{i}}{\partial x_{i}}=0,$ (2a)
$\frac{\partial u_{i}}{\partial t}+u_{j}\frac{\partial u_{i}}{\partial x_{j}}=-\frac{\partial p}{\partial x_{i}}+\nu\frac{\partial^{2}u_{i}}{\partial x_{j}\partial x_{j}},$ (2b)
where $p$ represents the kinematic pressure (i.e., static pressure divided by the fluid density) and $\bm{u}=[u,v,w]^{T}$ is the velocity.
To describe the flow at the macroscale, governing equations formulated in terms of averaged flow properties are necessary. The initial work in this field was undertaken by darcy1856, who established a linear relationship between the pressure drop and the velocity through a porous bed, under the creeping flow regime. When inertial effects become significant, the relationship between pressure gradient and velocity is no longer linear. A corrected equation accounting for fluids inertia by introducing a term proportional to a square of velocity was presented by Forchheimer1901 and is known as the Darcy–-Forchheimer equation for the pressure drop.
Although the DF model was originally postulated as empirical correlations, it has been successfully rederived by upscaling the equations valid at the pore scale (Whitaker1986; whitaker1996). Most of these derivations fundamentally rely on some form of averaging technique, such as Volume Averaging Method (VAM) (battiato2019). In the framework of VAM, an intrinsic average of a flow variable inside a porous medium is defined as
$\langle\psi\rangle=\frac{1}{V_{f}}\int\limits_{V_{f}}\psi\text{d}\,V,$ (3)
and is connected to a superficial average ${\langle\psi\rangle}_{s}$ by means of the Dupuit’s relation ${\langle\psi\rangle}_{s}=\phi{\langle\psi\rangle}$. A uniform flow in the vertical direction (along the $z$ coordinate axis) in the packed bed can be described as (woudberg2020):
$\frac{\partial\langle p\rangle}{\partial z}=-\frac{\nu}{K_{\mathrm{eff}}}\langle w\rangle_{s}\approx-\frac{\nu}{K}\left(1+C_{\mathrm{F}}\frac{\phi\sqrt{K}}{B}\mathrm{Re}_{\mathrm{p}}\right)\langle w\rangle_{s},$ (4)
where, $K_{\mathrm{eff}}=K_{\mathrm{eff}}(\mathrm{Re}_{\mathrm{p}})$ is the effective permeability of a porous material which accounts for both viscous and inertial effects, and is, therefore, a function of the Reynolds number. It can be approximated with the DF drag model formulated in terms of the isotropic permeability $K$ and the Forchheimer (or inertial) coefficient $C_{\mathrm{F}}$. The values of $K$ and $C_{\mathrm{F}}$ can be determined either experimentally, computed from simulations or estimated by using semi-empirical correlations (e.g., Lenci2022; woudberg2020; liu2024a). Alternatively, the effective permeability can be also expressed as:
$\frac{K_{\mathrm{eff}}}{B^{2}}\approx\frac{\mathrm{Da}}{1+C_{\mathrm{F}}\phi\sqrt{\mathrm{Da}}\mathrm{Re}_{\mathrm{p}}}.$ (5)
To compare results between different geometries, it is useful to express eq. 4 in terms on non-dimensional parameters. The pressure gradient can be normalized by the characteristic dimension of the studied geometry—taken here as BB—and the square of the superficial velocity, forming the friction factor (Dullien1975):
$f_{\mathrm{p}}=-\frac{B}{\langle w\rangle_{s}^{2}}\frac{\partial\langle p\rangle}{\partial z}.$ (6)
The permeability has the dimensions of length-squared leading to the definition of the Darcy number as $\mathrm{Da}=K/B^{2}$ and the following form of eq. 4
$f_{\mathrm{p}}\approx\frac{1+C_{\mathrm{F}}\phi\sqrt{\mathrm{Da}}\mathrm{Re}_{\mathrm{p}}}{\phi\mathrm{Re}_{\mathrm{p}}\mathrm{Da}}.$ (7)
4 Numerical Model
All simulations were performed using OpenFOAM-12 (greenshields2024), an open-source finite volume-based Computational Fluid Dynamics (CFD) software. Both steady-state and transient solvers were used, employing the SIMPLE and PISO pressure-velocity coupling methods, respectively (moukalled2016). To determine the appropriate solution method for each geometry and flow regime, both algorithms were tested to compute the flow at $\mathrm{Re}_{\mathrm{p}}=50$ in each geometry. The results showed no significant differences in the microscopic and macroscopic flow features, such as the pressure gradient and the intrinsic average of velocity indicating that the steady-state solver is an appropriate choice for this Reynolds number. Expecting the onset of unsteady flow for higher Reynolds numbers, all simulations with $\mathrm{Re}_{\mathrm{p}}>50$ were performed using the unsteady solver, otherwise a steady-state solver was used.
Table 1: The information regarding each of the considered geometries and computational meshes. The table contains the number of cells per layer ($N_{\mathrm{C}}$) and the number of layers ($N_{\mathrm{L}}$) for each of the considered rotation angles $\alpha$.
$\alpha$ $0^{\circ}$ $5^{\circ}$ $10^{\circ}$ $15^{\circ}$ $20^{\circ}$
$N_{\mathrm{L}}$ 1 36 18 12 9
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1 1.46 1.33 1.22 1.46
$\alpha$ $25^{\circ}$ $30^{\circ}$ $35^{\circ}$ $40^{\circ}$ $45^{\circ}$
$N_{\mathrm{L}}$ 36 6 36 9 4
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.21 1.27 1.30 1.27 1.24
$\alpha$ $0^{\circ} 55$$5^{\circ}$ $60^{\circ}$ $65^{\circ}$ $70^{\circ}$
$N_{\mathrm{L}}$ 18 36 3 36 18
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.22 1.24 1.26 1.28 1.27
$\alpha$ $75^{\circ}$ $80^{\circ}$ $85^{\circ}$ $90^{\circ}$
$N_{\mathrm{L}}$ 12 9 36 2
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.27 1.26 1.25 1.26
The residual convergence tolerance was set to $1\times 10^{-13}$ for pressure and velocity equations in steady-state simulations to ensure proper convergence of velocity gradients and accurate computation pressure drop over the packed bed geometry. The maximum number of correctors for mesh non-orthogonality was set to 3.
The unsteady simulations employed a less strict convergence tolerance equal to $1\times 10^{-8}$. The Courant number was kept below $0.8$. The number of PISO corrector steps was set to 4, with one additional loop for mesh non-orthogonality.
To force the flow, the pressure gradient momentum source was adopted, with the value calibrated during runtime, ensuring desired the bulk velocity and Reynolds number. Second-order accurate spatial discretization schemes using linear interpolation were used for both convective and diffusive terms. For time integration a second-order implicit backward scheme was adopted.
The native geometric–algebraic multigrid (GAMG) solver was predominantly used for the pressure equation in both steady and unsteady simulations. However, for rotation angles of $5^{\circ}$, $10^{\circ}$, $15^{\circ}$ and $20^{\circ}$, at high $\mathrm{Re}_{\mathrm{p}}$, the preconditioned conjugate gradient (PCG) method was used instead due to convergence difficulties. The PBiCGStab method with the DILU preconditioner was applied to the velocity equation in all cases.
Each module was meshed separately, resulting in the grids for each geometry consisting of a number of separate regions, which were connected through a non-conformal coupling (NCC) mesh interface (greenshields2024), which joins the overlapping cells and imposes wall boundary conditions otherwise. Side walls of the module were treated as no-slip boundaries with standard zero Neumann boundary conditions for pressure. The top and bottom boundaries of each geometry were also treated using NCC interface, imposing periodic boundary conditions to replicate an infinitely repeating porous structure.
An unstructured mesh was employed for all cases, generated using Gmsh (geuzaine2009), with prism cells near the corners of smaller void spaces.
Figure 3: (a) Schematic representation of the simulation domain for the case with a module angle of $50\text{\,}\mathrm{\SIUnitSymbolDegree}$. Part of the domain is clipped for clarity. The first zoomed-in view illustrates the orientation of the mesh, while the second highlights the mesh structure between adjacent layers. (b) Detailed view of the refined mesh at the same angle
Sensitivity of the results to the chosen resolution was evaluated using three different meshes without refinement (see appendix A for details). For the steady-state simulations, a medium-resolution mesh comprising approximately 1 million cells per module was selected. Depending on the rotation angle, this configuration produced a maximum non-orthogonality of $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ (with an average of $9.55\text{\,}\mathrm{\SIUnitSymbolDegree}$) and a maximum skewness of 0.84. As an example, a mesh with a rotation angle of $50\text{\,}\mathrm{\SIUnitSymbolDegree}$ is shown in fig. 3.
To accurately capture the unsteady flow, for transient simulations, the meshes were refined near the overlapping parts of the mesh interfaces between the regions. The resulting increase in cell count was geometry dependent: for the $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ case, this amounted to approximately 20%, whereas for the $5\text{\,}\mathrm{\SIUnitSymbolDegree}$ case, the increase was around 50% due to the larger contact surface. After refinement, the maximum non-orthogonality increased to about $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ (average of $12.28\text{\,}\mathrm{\SIUnitSymbolDegree}$), and the maximum skewness 1.34. A representative refined mesh for $50^{\circ}$ angle case is shown in fig. 3. Table 1 presents the number of cells for each rotated angle after the refinement.
Table 2: Normalized simulations times for different Reynolds numbers. Here, $T_{B}=B/\langle w\rangle$ denotes the characteristic time scale, $T_{\mathrm{i}}$ represents the start-up time, and $T_{\mathrm{avg}}$ the average time.
$\mathrm{Re}_{\mathrm{p}}$ $T_{\mathrm{i}}/T_{B}$ $T_{\mathrm{avg}}/T_{B}$
100 20 10
150 30 15
200 40 20
For transient simulations, the averaging time was defined such that the first-order quantities for each Reynolds number reached a statistically steady-state at randomly selected probe locations within the simulation domain. The times are listed in table 2.
4.1 Periodic Averaging & Data Reduction
Figure 4: Computational mesh of the $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ simulation domain, representing periodically repeated module used for averaging
For each rotation angle $\alpha$, the numerical domain consists of identical modules, repeated several times as described in section 2. The flow field in the void spaces of each module should be identical when viewed in a reference frame rotated with the module. This allows to further boost the statistical convergence of time-averaged data by performing a periodic average, or averaging the flow field over each module in the considered geometry.
This is achieved by first identifying the corresponding cells and faces, as illustrated in fig. 4. Scalar fields can then be directly averaged by summing the affiliated field values at these cell/faces and dividing by the total number of modules. In the case of vector and tensor fields, before averaging, the field has to be first rotated appropriately for each module. The average values can be distributed to the cells/faces corresponding to each other (applying the reverse transformation for the vector and tensor fields). Both the velocity and pressure fields are averaged this way in this study.
Finally, since the periodically averaged fields are identical across layers (without considering the rotation), only one representative module needs to be considered for analysis and data archiving. The boundary values of the representative module were calculated by interpolating between adjacent cells. As an example of the storage efficiency achieved, for $\alpha=$5\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=200$, the simulation comprising 36 layers originally required about 15 GB of storage, which was reduced to approximately 135 MB after averaging. We would like to remark that such method cannot be used for archiving of snapshots of unsteady velocity or pressure, and we employ it only for time-averaged data or out of steady-state simulations. Unless otherwise stated, the results presented in the following sections are based on periodically averaged mean values.
5 Results and Discussion
5.1 Validation
Figure 5: Comparison of normalized streamwise velocity fields $\overline{w}/\langle w\rangle$ obtained from PIV measurements and CFD simulations for two Reynolds numbers, $\mathrm{Re_{p}}=100$ and $200$. The upper figure (a) corresponds to P1, and the lower figure (b) corresponds to P3 within the packed bed module.
The numerical setup was validated by comparing results for $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=100,200$ against experiments conducted by christin2025. We compared the mean flow fields at two planes located between the bars (P1 and P3), as shown in fig. 1. The experimental configuration (referred to as RC2) consists of 18 modules and an outlet section, with measurements available at layers 13th–17th and above the bed. Because the velocity at the 13th layer is least affected by outlet effects, it was selected for comparison with the fully developed, periodic flow of the simulation.
The flow field at P1 and P3 is compared in fig. 5, plotted in a $(\xi,z)$ coordinate system (horizontal and vertical coordinates, respectively). The simulation captures the experimental flow features with satisfactory accuracy, particularly at the midspan region. Both datasets indicate that the flow field behavior is mainly determined by the geometry of the void space, especially by the locations of the inlets and outlets. The accelerated fluid near these narrow connections between the modules generates several recirculation regions attached to the bottom and top walls.
While the midspan velocity distributions show strong agreement in both flow direction and normalized velocity, deviations are observed at the far left and right sides of the domain. These discrepancies are attributed to the experimental setup: the RC2 geometry required modification to allow for optical access. For a comprehensive comparative analysis of the numerical and experimental results, the reader is referred to sadowski2025.
Figure 6: Normalized spanwise velocity profiles ($\overline{w}/\langle w\rangle$) from PIV measurements and CFD simulations. Profiles are shown at three vertical positions (z/B=12.25, 12.50, 12.75z/B=12.25,\,12.50,\,12.75) within the packed bed for $\mathrm{Re_{p}}=100$ and $200$.
Figure 6 presents the spanwise velocity profiles normalized by the intrinsic velocity at three $z$-positions of the 13th layer for $\mathrm{Re}_{\mathrm{p}}=100$ and $200$.
At $\mathrm{Re}_{\mathrm{p}}=100$, the simulations reproduce the characteristic flow features with high fidelity, capturing both the location and magnitude of the peak velocities. This quantitative agreement is confirmed by the root mean square deviation (RMSD) analysis, which yielded its lowest value (0.333) at this Reynolds number at location P3. However, three notable deviations are observed. First, minor discrepancies appear near the lateral boundaries, particularly at z=12.75mmz=12.75\ $\mathrm${mm}. Second, a boundary-layer jet is clearly resolved at P3 in the CFD data but is essentially absent in the PIV measurements; both are likely due to the geometric differences in the experimental setup discussed previously. Third, at the far left and right sides, P3 exhibits a jet-like flow while P1 does not (a trend also observed at $\mathrm{Re}_{\mathrm{p}}=200$). This is attributable to the location of P3, which is closer to the central axis, at this orientation.
At $\mathrm{Re}_{\mathrm{p}}=200$, the flow is unsteady, increased inertial effects lead to significantly stronger vortices and the emergence of distinct local velocity fluctuations. Reflecting this increased physical complexity, the RMSD at P1 is 1.009. Although this represents the maximum deviation across the simulated cases, it remains within acceptable limits considering the transient nature of the flow, and the overall agreement between the numerical and experimental results remains robust.
5.2 Overview of the Flow Fields
Figure 7: Normalized mean velocity magnitude without periodic averaging at $\mathrm{Re}_{\mathrm{p}}=200$ for (a) $20\text{\,}\mathrm{\SIUnitSymbolDegree}$ and (b) $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ geometries, visualized on the plane at position P3 (see fig. 1). The flow direction is along the +z+z axis.
We provide an overview of the flow field under different geometric configurations and flow conditions. Figure 7 presents the mean velocity field for two different angles ($\alpha=$20\text{\,}\mathrm{\SIUnitSymbolDegree}$,$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) at the highest Reynolds number ($\mathrm{Re}_{\mathrm{p}}=200$), on the slices through the void spaces in each module located at the position P3 (see fig. 1). The strong similarities observed between successive layers highlight the periodic nature of the flow, thereby justifying the use of periodically averaged quantities in the following sections.
For the rotation angle of $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, visualized in fig. 7, the areas of high velocity field are confined to the central section of the void spaces. This is a direct result of a channel-like nature of the geometries with small rotation angle. Each void space is characterized by a strong streamwise connectivity (closely aligned and single inlet and outlet) inducing preferential transport of momentum along the centre. Moreover, a weak recirculating flow can be observed in the “cavity”like regions of the void spaces located at the sides of the strong core flow. This is consistent with the observations of Zhang2025, that in the inertial regime of laminar flow, recirculation zones emerge near the walls when the flow passes through pore-throat structures.
[…]
Figure 8: Normalized mean velocity magnitude without periodic averaging in a packed bed with a rotation angle of $60\text{\,}\mathrm{\SIUnitSymbolDegree}$. Slices are taken from the middle of each module: (a) at $\mathrm{Re}_{\mathrm{p}}=1$, (b) at $\mathrm{Re}_{\mathrm{p}}=200$.
After examining the influence of the rotation angle on the flow structure, the effect of the Reynolds number on the flow regime was considered. Figure 8 shows the normalized mean velocity magnitude for a geometry with $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at two different Reynolds numbers. The slices shown in the fig. 8 are extracted from the mid-planes of three successive modules, located at $z=5$, $15$, and $25\text{\,}\mathrm{m}\mathrm{m}$, respectively.
At $\mathrm{Re}_{\mathrm{p}}=1$ (fig. 8), the flow is fully laminar and stationary with the pressure drop primarily governed by Darcy’s law. The fluid passes smoothly through the void spaces between layers, where the normalized velocity magnitude ranges from 0 to 2.5. The streamlines follow smooth slightly curved trajectories directed upward through the bed. The velocity distribution within the void space exhibits a low-speed, upward-directed flow that accelerates toward the center.
At $\mathrm{Re}_{\mathrm{p}}=200$ (fig. 8), the flow is in the inertia-dominated regime. In contrast to $\mathrm{Re}_{\mathrm{p}}=1$ (fig. 8), the streamlines here exhibit significant tortuosity. This complexity arises because the increased inertial forces prevent the fluid from following the curvature of the solid bars,leading to the formation of low-velocity wake regions and distinct recirculation zones within the void spaces. The flow develops strong transverse velocity components, which drive significant deviations from the primary streamwise direction.
5.3 Friction Factor
Figure 9: Friction factor ($f_{\mathrm{p}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different rotation angles
Figure 9 shows the relation of the friction factor as a function of Reynolds number for various angles. Each line represents the friction factor calculated from eq. 6. The color gradient indicates the rotation angle of each geometry from $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$.
In the viscous regime, all curves follow a hyperbolic trend with respect to $\mathrm{Re}_{\mathrm{p}}$, which appears as a straight line in the log–log scale. As the angle increases from $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the friction factor increases and reaches its maximum at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Beyond this angle, the friction factor decreases gradually, and further changes in rotation angle no longer introduce a noticeable deviation. In other words, within this regime, all curves nearly collapse onto one another, indicating that the influence of the rotation angle becomes weak.
The specific alignment of the square bars at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ creates a severe constriction of the flow path. These narrowing void spaces increase the local shear stress, resulting in the maximum friction factor observed in the viscous regime.
As the Reynolds number increases, deviations from the linear trend become apparent, signaling the onset of inertial effects and the transition toward the Forchheimer regime. Additionally, the angle yielding the maximum friction factor is found to shift with the Reynolds number. For instance, at $\mathrm{Re}_{\mathrm{p}}=200$, the maximum friction factor occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, in contrast to the peak at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ observed in the viscous regime.
At $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the flow shows behavior similar to contraction–expansion sequences in ducts. As the Reynolds number increases, the acceleration into void spaces creates local jets, leading to flow separation and recirculation downstream of these constrictions. Consequently, form drag becomes the dominant loss mechanism, resulting in the highest observed friction factor
One of the most notable trends is the nearly linear behavior for $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, where perfect orthogonality between modules allows the flow to develop with minimal obstruction, resulting in reduced flow resistance. Therefore, the $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration can also be classified as a channel-like geometry.
Figure 10 presents a comparison between the computed friction factor ($f_{\mathrm{p}}$), obtained from eq. 6, and the fitted non-dimensional friction factor ($f_{\mathrm{fitted}}$). The latter was evaluated using eq. 7, incorporating the calculated Darcy number and Forchheimer coefficient. As shown in fig. 10, the fitted curves closely follow the computed data over the selected range of Reynolds numbers. The agreement is particularly strong in the viscous regime, while slight deviations appear at higher Reynolds numbers, suggesting that additional simulations at higher $\mathrm{Re}_{\mathrm{p}}$ may be required to better capture the inertial effects.
5.4 Fitting Model Parameters
Figure 10: Comparison between the computed friction factor ($f_{\mathrm{p}}$) and the fitted friction factor ($f_{\mathrm{fitted}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ number for different angles. Each curve corresponds to a distinct angle ranging from $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, with the indicated scaling factors applied to separate the curves vertically for clarity.
Figure 11: Variation of (a) Darcy number and (b) Forchheimer coefficient w.r.t. the rotation angle ($\alpha$)
The determination of the $\mathrm{Da}$ and $C_{\mathrm{F}}$ numbers was carried out in two steps using a least-squares fitting optimization procedure. First, the Darcy number was obtained for $\mathrm{Re}\leq 1$ by neglecting inertial effects (i.e. assuming $C_{\mathrm{F}}=0$) and using eq. 5. In the second step, the $C_{\mathrm{F}}$ was evaluated by fitting the eq. 5 over the entire Reynolds number range while keeping the Darcy number fixed.
The Darcy number shows a strong dependence on the rotation angle, demonstrating two distinct behaviours depending on the geometries as shown in fig. 11. For $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structures are aligned in a channel-like form, which promotes a less perturbed flow and results in high permeability. Beyond this angle, the geometry becomes multiply connected, leading to a sharp drop in permeability. For $\alpha>$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the Darcy number exhibits a sinusoidal-like oscillation with increasing angle, showing reduced sensitivity to the rotation angle. Interestingly, although $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ produces a channel-like flow, its permeability does not follow the behaviour typically we observe in similar geometries. Instead, the Darcy number (or permeability) at $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ remains significantly lower than that observed at small rotation angles such as $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ or $5\text{\,}\mathrm{\SIUnitSymbolDegree}$. This discrepancy highlights the difficulty of predicting drag and permeability in such geometries.
[…]
We also report the Forchheimer coefficient in fig. 11 as a function of the rotation angle. It should be noted that the inertial regime was not fully developed within the current Reynolds number range; higher Reynolds numbers would be required to accurately fit the inertial contribution of the pressure drop. Therefore, the presented correlations can be considered only valid within the stated range of Reynolds number. Figure 11 indicates that the highest Forchheimer coefficient, $C_{\mathrm{f}}$, occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, which is consistent with the angle exhibiting the largest friction factor at $\mathrm{Re}_{\mathrm{p}}=200$.
5.5 Transition to Inertial Regime
Figure 12: Relative deviation of the friction factor from the $f_{\mathrm{p}}$. The circular markers with black outlines show the critical Reynolds number for each angle.
According to the classification proposed by Dybbs1984ANL, this transition typically occurs near $\mathrm{Re}_{\mathrm{p}}\approx 1$ for spherical particles in packed beds. The present geometry consists of square bars arranged at a constant rotation angle, leading to void space connectivity characteristics that differ significantly from those of a sphere-packed medium.
Hlushkou2006 suggested that if the inertial contribution exceeds $5\%$, the flow should no longer be considered in the Darcy regime. The same criterion is adopted in the present study. To identify the Reynolds number at which inertial effects become significant, the difference between the total friction factor and the linear friction factor ($f_{\mathrm{lin}}=1/(\phi\,\mathrm{Re}_{\mathrm{p}}\,\mathrm{Da})$) is computed and then divided by the total friction factor. We applied the transition criterion proposed by Hlushkou2006 to identify the onset of inertia-dominated flow. However, other criteria have also been reported in the literature, such as introduced by Zeng2006; bagci2014. In future work, different transition criteria may be used to provide a more comprehensive assessment of the transition to inertia-dominated flow.
We illustrate the transition from viscosity to inertia-dominated regime in fig. 12. The upper figure shows the relative friction factor for each angle. As expected, the deviation remains zero for $\mathrm{Re}_{\mathrm{p}}\leq 1$. With increasing $\mathrm{Re}_{\mathrm{p}}$, the inertial contribution represented by the Forchheimer term grows with the Reynolds number. Consequently, the pressure drop no longer follows the linear Darcy regime, and the deviation from Darcy’s law increases rapidly.
At $\mathrm{Re}_{\mathrm{p}}=200$, the rotation angles $\alpha\in\{0^{\circ},5^{\circ},\dots,25^{\circ}\}$ and $\{85^{\circ},90^{\circ}\}$ deviate notably from the behaviour observed at the remaining angles. For these other angles, the nonlinear component accounts for 80–90% of the total friction factor. The primary mechanism driving this distinct behaviour is the formation of channel-like flow paths. Although some of these angles (such as $25\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $85\text{\,}\mathrm{\SIUnitSymbolDegree}$) do not form fully developed channel-like geometries, they still generate localized preferential jet-like flow within the void space.
[…]
At $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the transition to the inertial regime occurs at approximately $\mathrm{Re}_{\mathrm{p}}\approx 5$, which is lower than for the other geometries. Potential factors contributing to this behavior include the sensitivity of the Darcy number evaluation for this orientation and the unique connectivity of the void space, which may promote early localized recirculation. A conclusive explanation would require a detailed flow field analysis which currently exceeds the scope of this study.
6 Conclusion
In this study, numerical simulations were performed using OpenFOAM-12 to investigate viscosity- and inertia-dominated flow through periodically arranged square-bar structures over a range of $\mathrm{Re}_{\mathrm{p}}=0.1\text{--}200$ with different angles ($0\text{\,}\mathrm{\SIUnitSymbolDegree}$$90\text{\,}\mathrm{\SIUnitSymbolDegree}$) at fixed porosity $\phi=0.332$. A total of 266 simulations were performed using a predominated hexahedral mesh, and selected cases were successfully validated against available PIV measurements obtained from the experimental configuration of christin2025. Some important conclusions can be drawn from the obtained results:
  1. (1)
    The rotation angle has a strong influence on the macroscopic flow characteristics. For $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure forms channel-like geometries, resulting in high effective permeability over the entire $\mathrm{Re}_{\mathrm{p}}$ range and low flow resistance at high $\mathrm{Re}_{\mathrm{p}}$. As the angle increases, the connectivity between successive layers becomes more complex, leading to the development of recirculation zones and increased frictional losses, resulting in heterogeneous velocity fields. We refer to these configurations as lattice-like geometries. At $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure again forms a mainly channel-like arrangement. In this orientation, the flow path becomes a less perturbed, and the disturbances observed at other rotation angles are reduced.
  2. (2)
    For the investigated Reynolds numbers, the onset of te inertial regime was defined as the critical Reynolds number, which was identified for each angle by checking when the inertial contribution to the friction factor exceeded 5%. For lattice-like geometries, the transition occurs at $\mathrm{Re}_{\mathrm{p}}\approx 7.5$, with the earliest onset at $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. When the rotation angle exceeds $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, the inertial regime begins at higher Reynolds numbers.
  3. (3)
    At $\mathrm{Re}_{\mathrm{p}}=0.1$, the highest friction factor was obtained at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas at $\mathrm{Re}_{\mathrm{p}}=200$, it occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. This demonstrates that the geometrical configuration exerts a strong influence on flow resistance, even at constant porosity.
  4. (4)
    It is important to note that the present numerical study was conducted at a fixed porosity to effectively highlight the impact of the inertial structure induced by the rotation angle. This influence of varying geometry is typically not included in standard permeability models, which are generally formulated as a function of porosity. Future work should consider modelling this geometric influence using approaches similar to those of (duplessis1988; VANWACHEM2024104815), validated against the present dataset.
Overall, the good agreement between simulations and experiments shows that the present configuration provides a reliable framework for analyzing flow through such periodic structures. Future work will extend the analysis to higher Reynolds numbers to examine the onset of unsteady and turbulent flow in more detail.
7 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8 Acknowledgments
The authors gratefully acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG) through No. SFB/ TRR287, Project No. 422037413.
Appendix A Effect of Mesh Resolution on the Numerical Results
Figure 13: Normalized spanwise-averaged velocity profiles at P3 (see fig. 1) for angles $30\text{\,}\mathrm{\SIUnitSymbolDegree}$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. Results are shown for three mesh resolutions (200K, 1M, and 4M cells)
Figure 14: Friction factor for three mesh sizes (200K, 1M, and 4M cells) at angles of $30^{\circ}$, $60^{\circ}$, and $90^{\circ}$.
Table 3: Friction factors used in for each angle and mesh. The table contains the number of cells per layer ($N_{\mathrm{C}}$)
$\alpha$ $N_{\mathrm{C}}$ $2\text{\times}{10}^{5}$ $1\text{\times}{10}^{6}$ $4\text{\times}{10}^{6}$
$30\text{\,}\mathrm{\SIUnitSymbolDegree}$ 67.70 70.27 72.87
$60\text{\,}\mathrm{\SIUnitSymbolDegree}$ 85.33 87.84 91.01
$90\text{\,}\mathrm{\SIUnitSymbolDegree}$ 16.36 14.53 14.47
A mesh sensitivity analysis to the results was performed using three unrefined meshes containing approximately $2\text{\times}{10}^{5}$, $1\text{\times}{10}^{6}$, and $4\text{\times}{10}^{6}$ cells per module at $\mathrm{Re}_{\mathrm{p}}=200$. No local refinement was applied in the flow region. Three modules with different rotation angle, namely $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, were considered. The objective of this analysis was to assess the influence of the overall mesh resolution on the solution accuracy and to determine an optimal mesh for the subsequent simulations.
The analysis was conducted for the case shown in fig. 13, focusing on the first module of the computational domain along the probe line located at point P3 (see fig. 1), positioned at the mid-line in the $z$-direction. Velocity data were extracted from statistically steady-state solutions without applying periodic averaging. Figure 13 presents the normalized spanwise-averaged velocity profiles. The coarsest mesh ($2\text{\times}{10}^{5}$ cells) shows a noticeable deviation from the finer ones, particularly in the central region of the $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ case. For $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the deviation becomes evident near the peak locations of the velocity profiles, where the $2\text{\times}{10}^{5}$ cell mesh results in less accurate results. As the mesh is refined, the profiles converge toward those of the finest mesh, indicating improved resolution of the flow field. For all investigated angles, the $1\text{\times}{10}^{6}$ and $4\text{\times}{10}^{6}$ meshes yield nearly identical profiles, suggesting that the flow features are well-resolved from the $1\text{\times}{10}^{6}$ cell mesh.
[…]
The mesh sensitivity analysis demonstrates that the $1\text{\times}{10}^{6}$ cell mesh provides reliable results while maintaining a reasonable computational cost.