### v3

# Modelling of pressure drop in periodic square-bar packed beds

Abstract
Understanding fluid flow through porous media with complex geometries is essential for improving the design and operation of packed-bed reactors. Most existing studies focus on spherical packings, having as a consequence that accurate models for irregular interstitial geometries are scarce. In this study, we numerically investigated the flow through a set of packed-bed geometries consisting of square bars stacked on top of each other and arranged in disk-shaped modules. Rotation of each module allows the generation of a variety of geometrical configurations at Reynolds numbers of up to 200 (based on the bar size). Simulations were carried out using the open-source solver OpenFOAM. Selected cases (e.g., $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\mathrm{Re}_{\mathrm{p}}=100,200$) were compared against Particle Image Velocimetry measurements. Results reveal that, based on the relative rotation angle, the realized geometries can be classified as channel-like ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and lattice-like ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), fundamentally altering the friction factor. Furthermore, the maximum friction factor obtained in the creeping regime occurred at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas in the inertial regime, this occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. The module-equivalent diameter, based on the angle-dependent wetted surface area, collapses the friction factor onto the Ergun correlation and yields good permeability predictions for the lattice-like geometries.
keywords:
Pressure drop , Packed bed , Friction factor , Permeability , Tortuosity , Non-spherical porous medium
journal: Particuology
[…]
[label1]organization=Ruhr University Bochum, addressline=Department of Mechanical Engineering, city=Bochum, postcode=44801, state=North Rhine-Westphalia, country=Germany
1 Introduction
A packed bed consists of a solid packing, typically a particle assembly confined within an enclosure, and passed by a fluid. Packed beds are widely used in petroleum, process, and chemical industry (seckendorff21). Typical examples are catalyst reactors (MiguelGarcia20), where residual methane undergoes total oxidation to produce carbon dioxide and water; catalytic converters for emission control, e.g. (Schnitzlein87), and thermal energy storage systems (RyanAnderson15), or as filters in environmental engineering (ChenWang19). The solid particles are often modelled as monodisperse spheres; in reality, however, they can span wide ranges of sizes and shapes, e.g., cylinders, polyhedra, or arbitrary non-polyhedral forms such as in (Afandizadeh2001; Schlipf2015).
The presence of the solid phase forces the fluid, in general, to adopt complex meandering paths as it permeates the packing, enhancing the heat and mass transfer and facilitating reactions (Dixon20). The gas flow in the interstitial space, which is the focus of this study, is influenced by both micro- and macroscopic parameters of the packing, e.g., particle size and shape (or pore size), bed porosity and permeability, and the operating conditions of the device characterized for example, by the Reynolds number (Kumar23; Abdulmohsin17). Therefore, a clear understanding of the impact of these factors on the flow features is crucial for designing efficient reactors (Elzubeir24).
Due to the complexity of these systems, detailed investigation, both numerical and experimental, covering all relevant scales is not viable. Therefore, numerical models of varying fidelity are generally used in the design process. For example, the whole device can be represented with a fast, one-dimensional model neglecting the geometrical complexity (niedermeier2018; Nash2017, e.g.,). When the problem has to be modelled in two or three dimensions, the flow field can be described at a macroscale level (i.e., using averaged flow properties) using porous medium models (collazo2012; battiato2019; Sadowski2023). This typically involves modelling the drag induced by the porous medium using closures (Whitaker1986; whitaker1996) such as Darcy–Forchheimer (DF). Complementary to these continuum approaches, particle movement and macroscopic parameters of the particle assembly can be tracked by the Discrete Element Method (Cundall1979; illanamahiques2023; Ma2022).
Higher accuracy of representation of the flow field is possible resolving the geometry of the packing, and ensuring a sufficiently accurate spatio-temporal discretisation to capture all relevant scales (Dixon20). Such approach—called Particle-Resolved Direct Numerical Simulation (PRDNS)—offers the most precise representation of both micro- and macroscale flow properties within packed beds. PRDNS can provide detailed data on interphase momentum, energy, and mass transfer allowing to develop, calibrate, and validate closure models for unresolved simulations (Tenneti2014; Sadowski2023b; Sadowski2023). However, such simulation requires high computational effort and ,hence, is typically limited to simple and laboratory-scale configurations (neeraj2023; sadowski24; sadowski2025).
A further current limitation of investigations of particle assemblies and packed beds is the simplification of the particle geometry, whereby particles are generally considered as spherical (e.g., zhu16; dentz22; storm24), even in approximations of industrial devices (neeraj2023; sadowski24), for example in a body-centered cubic (BCC) arrangement. Spherical particles offer practical advantages as the calculation of wall-distances can be enormously simplified and the definition of particle contact regions is relatively straightforward. This is more challenging for irregular shapes (Dixon20; jurtz2019) so that spherical particles allow for efficient contact detection and simulation of full-scale systems (illanamahiques2023a).
However, the higher complexity of the flow field patterns induced by heterogeneous, non-spherical packings can result in large errors in velocity predictions by models describing the flows in terms of averaged velocity (Moghaddam19). Moreover, roding2017 observed that the effective diffusivity in random packings of cuboids, spheres and ellipsoids is significantly influenced by the particle shape at identical porosity. This is troubling from the perspective of unresolved flow modelling, as most models express the drag or diffusivity in terms of macroscopic parameters and would be unable to capture such differences. Standard correlations for pressure drop, such as the Ergun equation (ergun1951) or Kozeny–Carman model (carman1937), were originally developed only for spherical particles. These and similar drag laws have, to some extent, been successfully generalized to other geometries (e.g., liu94; duplessis1988; duplessis2008) and adjusted to take into account the presence of other particle shapes, either by including the sphericity of the particles as a parameter of the model (woudberg2020; Li2011EffectiveDiameter), or by constructing a multiscale model accounting for average anisotropy of the packing based on microscale simulations (tobis2008; tobis2000). That said, to the best of our knowledge, the influence on the induced drag by the variation in packing geometry at a constant macroscopic geometrical properties (e.g, porosity) remains largely unquantified.
The geometry investigated in this study is based on the modular packed-bed reactor designed by christin2025, which consists of disk-shaped modules containing square bars that can be rotated relative to one another. This allows for the study of a wide range of interstitial configurations can be studied, while keeping the porosity, bed cross-section and the shape of modules strictly constant. Although this geometry does not represent a specific industrial packing, it provides a controlled framework for studying flow through complex interstitial geometries. Crucially, the availability of detailed Particle Image Velocimetry (PIV) measurements (christin2025) enables direct validation of the numerical results, which is rarely possible for non-spherical packed beds.
Therefore, this study explores how the variation of the interstitial geometry affects the induced drag and the permeability in a dense packed bed consisting of square bars. To quantify these effects, we carried out numerical simulations in a set of different geometries with the same porosity ($\phi=0.322$) at Reynolds numbers (based on bulk velocity and bar size) spanning the range between $0.1$ and $200$. Special attention was given to the influence of the geometry on the transition from viscosity- to inertia-dominated flow regimes. Moreover, to ensure the quality of the comprehensive database resulting from the 266 simulations, we carried out a detailed validation of the numerical setup against the PIV results of christin2025. Furthermore, a thorough mesh sensitivity analysis was performed to assess the viability of the numerical approach.
The remaining part of this work is organized as follows. In Section 2, the geometry of the packed bed and the associated flow conditions are described in detail. In Sections 3 and 4 the mathematical and numerical models used in the study, respectively, are described. Section 5 presents the results with particular emphasis on the non-dimensional parameters in porous media. Finally, we summarize our conclusion in Section 6.
Figure 1: (a) The laboratory-scale packed bed of christin2025. (b) The cross-section of the packed-bed geometry (christin2025) formed from six modules, each rotated by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ relative to the preceding one. (c) Schematic of the geometry of each module showing the circle enclosing the dodecagon defining the outer geometry of the slits. The streamwise direction is oriented perpendicular to the surface of the paper.
Figure 2: Visualizations of the simulation domains (i.e., volumes occupied by the fluid) for each of the studied rotation angles, ranging from $5^{\circ}$ to $90^{\circ}$. The streamwise direction is oriented vertically, with the flow moving from bottom to top.
2 Geometry and flow conditions
The geometry investigated in this work is based on a laboratory-scale modular packed-bed reactor designed by christin2025. The reactor, shown in Fig. 1, is a column consisting of identical modules and an outlet section which is not considered in this study. Each module is a disk with a thickness of $B=$10\text{\,}\mathrm{m}\mathrm{m}$\$, with $5\text{\,}\mathrm{m}\mathrm{m}$ wide slits cut through, forming prismatic bars with a square cross-section ($B\times B$) within each module. The corresponding dimensions are shown in Fig. 1. The geometry of the void space in each module is defined so that the side walls of the slits are lying on a regular dodecagon inscribed in a circle with a diameter $D=$62.12\text{\,}\mathrm{m}\mathrm{m}$\$, which is concentric with the modules outer geometry. The area of this circle defines the bed’s cross-sectional area, leading to a theoretical porosity of $\phi=0.322$.
The disks are stacked vertically and rotated relative to one another, creating a complex void-space geometry between adjacent layers. An example of the resulting geometry, with a rotation angle of $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ between layers, is shown in Fig. 1.
This geometry was designed to allow for various geometries of interstitial space: the rotation angle, height of the modules or the size of the bars can be adjusted independently. While not directly tied to a specific industrial application, the use of bars in each module enables spatio-temporally resolved optical measurements in each configuration (christin2025), unlike typically encountered, random non-spherical packing. This last property makes it valuable for studying flow characteristics via simulation approaches, as the available measurement data allows for the detailed validation at the pore scale.
By rotating each module by a fixed angle, different interstitial configurations and flow characteristics in the bed can be investigated. In the present study, $19$ rotation angles $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$,$5\text{\,}\mathrm{\SIUnitSymbolDegree}$,\dotsc,$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ and, thereby, 19 geometries, are considered. In order to accurately represent the macroscopic flow properties of each of packing, it is vital that a periodic geometry is constructed (guibert2016; scandelli2022), otherwise a blockage effects would influence the computed values of pressure drop. For this reason, for each rotation angle, such a number of layers is chosen, which results in a periodically repeating geometry. The number of layers corresponding to each rotation angle is provided in Table 1.
Figure 2 shows the simulation domains corresponding to each angle. The studied geometries can be classified by specific surface area $a_{v}$ defined as the ratio of the area of the wetted surface to the total volume of the packed bed, and the ratio of the diameter of the bed to the equivalent particle diameter $D/d_{\mathrm{eq}}$, both plotted in Fig. 3 (see Section 3 for detailed discussion of evaluating both parameters for a given geometry).
At small angles the bars in adjacent modules are nearly aligned, creating large areas of contact between the bars and yield the lowest values of $a_{v}$. As $\alpha$ increases, more surface is exposed, so that $a_{v}B$ rises to about $2.0$ near $\alpha=25^{\circ}$. For $\alpha\geq 15^{\circ}$$a_{v}B$ oscillates with $\alpha$, reflecting the periodically changing contact pattern between successive layers. Since $d_{\mathrm{eq}}\propto 1/a_{v}$ (Eqs. 17 and 8), $D/d_{\mathrm{eq}}$ follows the same trend, ranging from approximately $1.9$ at $\alpha=0^{\circ}$ to 3.13.13.33.3 for $\alpha\geq 15^{\circ}$.
For the current geometry, we picked $D/d_{\mathrm{eq}}\approx 3$ to marks the boundary between the channel-like ($D/d_{\mathrm{eq}}<3$) and lattice-like ($D/d_{\mathrm{eq}}>3$) geometries, with the transition occurring between $\alpha=10^{\circ}$ and $15^{\circ}$. Therefore, for small rotation angles ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), the arrangement resembles a channel-like structure, where aligned void spaces create curved flow passages with minimal obstruction. As $\alpha$ increases beyond $10\text{\,}\mathrm{\SIUnitSymbolDegree}$, the void spaces become increasingly interconnected and the geometry transitions to a lattice-like structure with multiply connected flow paths. The $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration, with $D/d_{\mathrm{eq}}=3.23$ and $a_{v}=0.193~\mathrm{mm^{-1}}$, is classified within the lattice-like group based on these geometric parameters (see Section 5).
Figure 3: Non-dimensional specific surface area $a_{v}B$ and bed-to-particle diameter ratio $D/d_{\mathrm{eq}}$ as a function of the rotation angle $\alpha$
These values of $D/d_{\mathrm{eq}}$ place the RC2 geometry in the low bed-to-particle diameter ratio regime ($D/d<5$) (yang2016), comparable to the random packings of spheres and cylinders studied by Moghaddam19 (N=2.29N=2.296.16.1) and papkov2024 ($D/d=2.9$ and $4.8$).
The fluid moves through the column in the axial direction, with the flow conditions characterized by the pore Reynolds number The pore Reynolds number is defined as
$\mathrm{Re}_{\mathrm{p}}=\frac{\langle w\rangle B}{\nu},$ (1)
where $B$ corresponds to the bar size, $\langle w\rangle$ is the intrinsic average of the streamwise velocity, and $\nu$ is the kinematic viscosity. (In the experimental setup of (christin2025), the intrinsic average of velocity is computed directly from the volumetric flow rate forced through the packed bed.)
Depending on the application, the flows in packed beds and porous materials can be characterized by a wide range of Reynolds numbers (wood2020; Dullien1975), requiring the accurate description of both the viscous and inertial effects, to properly model the flow behaviour. Therefore, in the present paper we focus on a Range of Reynolds numbers spanning $0.1$ to $200$.
According to the flow regime classification of (Dybbs1984ANL), the transition to turbulence is typically reported at the pore-size Reynolds numbers in the range $\mathrm{Re}_{p}\approx 300$$500$ for packed beds of spheres. (wood2020), in their review of turbulent flows in porous media, confirmed that these thresholds are broadly consistent across different packing configurations, although the exact transition depends on the specific geometry, and well-defined turbulence can be established above $\mathrm{Re}_{p}\sim 750$. Since the present study is limited to $Re_{p}\leq 200$, all considered flow conditions can be expected to be laminar, with, for higher Reynolds numbers, strong inertial effects and unsteady behaviour which can exhibit intermittency.
3 Mathematical model
An incompressible flow of a Newtonian fluid with constant and homogeneous density (i.e., neglecting variations due to mixing, reaction, or thermal effects) inside the considered packed bed can be described at the pore level by the Navier–Stokes and continuity equations:
$\frac{\partial u_{i}}{\partial x_{i}}=0,$ (2a)
$\frac{\partial u_{i}}{\partial t}+u_{j}\frac{\partial u_{i}}{\partial x_{j}}=-\frac{\partial p}{\partial x_{i}}+\nu\frac{\partial^{2}u_{i}}{\partial x_{j}\partial x_{j}},$ (2b)
where $p$ represents the kinematic pressure (i.e., static pressure divided by the fluid density) and $\bm{u}=[u,v,w]^{T}$ is the velocity.
To describe the flow at the macroscale, governing equations formulated in terms of averaged flow properties are necessary. The initial work in this field was undertaken by darcy1856, who established a linear relationship between the pressure drop and the velocity through a porous bed, under the creeping flow regime. When inertial effects become significant, the relationship between pressure gradient and velocity is no longer linear. A corrected equation accounting for fluid inertia by introducing a term proportional to the square of the velocity was presented by Forchheimer1901 and is known as the Darcy–Forchheimer equation for the pressure drop.
Although the DF model was originally postulated as empirical correlations, it has been successfully rederived by upscaling the equations valid at the pore scale (Whitaker1986; whitaker1996). Most of these derivations fundamentally rely on some form of averaging technique, such as the Volume Averaging Method (VAM) (battiato2019). In the framework of VAM, an intrinsic average of a flow variable inside a porous medium is defined as
$\langle\psi\rangle=\frac{1}{V_{f}}\int\limits_{V_{f}}\psi\text{d}\,V,$ (3)
and is connected to a superficial average ${\langle\psi\rangle}_{s}$ by means of the Dupuit’s relation ${\langle\psi\rangle}_{s}=\phi{\langle\psi\rangle}$. A uniform flow in the vertical direction (along the $z$ coordinate axis) in the packed bed can be described as (woudberg2020):
$\frac{\partial\langle p\rangle}{\partial z}=-\frac{\nu}{K_{\mathrm{eff}}}\langle w\rangle_{s}\approx-\frac{\nu}{K}\left(1+C_{\mathrm{F}}\frac{\phi\sqrt{K}}{B}\mathrm{Re}_{\mathrm{p}}\right)\langle w\rangle_{s},$ (4)
where $K_{\mathrm{eff}}=K_{\mathrm{eff}}(\mathrm{Re}_{\mathrm{p}})$ is the effective permeability of a porous material which accounts for both viscous and inertial effects, and is, therefore, a function of the Reynolds number. It can be approximated with the DF drag model formulated in terms of the isotropic permeability $K$, which is a limit of $K_{\mathrm{eff}}$ at zero Reynolds number, and the Forchheimer (or inertial) coefficient $C_{\mathrm{F}}$. The values of $K$ and $C_{\mathrm{F}}$ can be determined experimentally, computed from simulations, or estimated using semi-empirical correlations (e.g., Lenci2022; woudberg2020; liu2024a). The effective permeability can be non-dimensionalised by selecting a characteristic length scale of the studied problem, here $B$, and forming the Darcy number $\mathrm{Da}_{\mathrm{(eff)}}=K_{\mathrm{(eff)}}/B^{2}$:
$\frac{K_{\mathrm{eff}}}{B^{2}}\approx\frac{\mathrm{Da}}{1+C_{\mathrm{F}}\phi\sqrt{\mathrm{Da}}\mathrm{Re}_{\mathrm{p}}}.$ (5)
It is also useful to express the pressure drop in terms of a friction factor, as commonly done in the packed-bed literature (Dullien1975). The pressure gradient can be normalized by the characteristic dimension of the studied geometry, taken here again as $B$, and the square of the superficial velocity:
$f_{\mathrm{p}}=-\frac{B}{\langle w\rangle_{s}^{2}}\frac{\partial\langle p\rangle}{\partial z},$ (6)
which combined with Eq. 5 gives:
$f_{\mathrm{p}}\approx\frac{1+C_{\mathrm{F}}\phi\sqrt{\mathrm{Da}}\mathrm{Re}_{\mathrm{p}}}{\phi\mathrm{Re}_{\mathrm{p}}\mathrm{Da}}.$ (7)
Importantly, the choice of characteristic dimension is not unique and other definitions of the friction factor have been proposed. A common approach uses the specific surface area $a_{v}=A_{p}/V_{p}$ of a particle to define the Sauter-mean diameter
$d_{\mathrm{sd}}=\frac{6}{a_{v}}=\frac{6\,V_{p}}{A_{p}}\,.$ (8)
where $V_{p}$ and $A_{p}$ are the volume and surface area of the particle, respectively. The Sauter-mean diameter does not account for particle non-sphericity. To incorporate shape effects, the sphericity $\psi$ (Wadell1935) is introduced as the ratio of the surface area of a volume-equivalent sphere to that of the actual particle:
$\psi=\frac{\pi^{1/3}\left(6\,V_{\mathrm{p}}\right)^{2/3}}{A_{\mathrm{p}}}\,.$ (9)
Li2011EffectiveDiameter proposed the equivalent particle diameter, defined as
$d_{\mathrm{eq}}=\psi\,d_{\mathrm{sd}}\,$, (10)
which reduces to $d_{\mathrm{sd}}$ for spherical particles where $\psi=1$. Using such length scale, the ergun1951 equation can be used to model the drag in packed beds of non-spherical particles (li2011), which is often expressed in the form of the following friction factor:
$f_{\mathrm{Erg}}=-\frac{\partial\langle p\rangle}{\partial z}\frac{d_{\mathrm{eq}}}{\langle w\rangle_{s}^{2}}=\frac{1-\phi}{\phi^{3}}\left(\frac{150}{\mathrm{Re}_{\mathrm{Erg}}}+1.75\right),$ (11)
with the Reynolds number defined as
$\mathrm{Re}_{\mathrm{Erg}}=\frac{\langle w\rangle d_{\mathrm{eq}}}{\nu}\frac{\phi}{1-\phi}.$ (12)
Equation 11 can be recast in the form equivalent to Eq. 4 with the Blake–Kozeny model for permeability
$K^{\mathrm{Erg}}=\frac{d_{\mathrm{eq}}^{2}}{150}\frac{\phi^{3}}{(1-\phi)^{2}}$ (13)
and the Forchheimer coefficient expressed as:
$C^{\mathrm{Erg}}_{\mathrm{F}}=\frac{1.75\,(1-\phi)\,\sqrt{K/B^{2}}}{\phi^{3}}.$ (14)
The definitions of the friction factor fundamentally depend on how the characteristic dimension is computed, as it influences both the definition of the Reynolds number and the coefficients of drag. Because the RC2 module contains interconnected bars rather than loose particles, two definitions of characteristic diameters can be considered: one based on the module geometry as a whole, and one based on a single representative bar. These are evaluated in the following section.
3.1 Geometric parameters for the RC2 configuration
In the present work, the specific surface area is primarily defined as
$a_{v}=\frac{A_{\mathrm{wetted}}}{V_{\mathrm{total}}}\,$, (15)
where $A_{\mathrm{wetted}}$ is the total wetted surface area of the module and $V_{\mathrm{total}}$ is the total module volume (comprising both fluid and solid regions). Because the bar-to-bar contact regions change with the rotation angle $\alpha$, so does the wetted surface exposed to the fluid, and consequently both $a_{v}$ and the bed-to-particle diameter ratio $D/d_{\mathrm{eq}}$ vary with $\alpha$ (Fig. 3). The equivalent diameter $d_{\mathrm{eq}}$ is derived from the module as a whole. The module-based sphericity is
$\psi_{\mathrm{eq}}=\frac{\pi^{1/3}\,\bigl(6\,V_{\mathrm{p}}\bigr)^{2/3}}{A_{\mathrm{wetted}}}\,$, (16)
The Sauter-mean diameter follows directly from Eq. 8 and likewise depends on $\alpha$. The module-based equivalent diameter then follows from Eqs. 10 and 8 as
$d_{\mathrm{eq}}=\psi_{\mathrm{eq}}\,d_{\mathrm{sd}}^{\mathrm{eq}}\,.$ (17)
One can also propose a characteristic length scale by establishing a representative particle and use it to compute geometric properties: based on a single representative square bar, we define the single-bar diameter $d_{\mathrm{sb}}$. The simulation domain per module contains four square bars arranged in two symmetrically positioned pairs that share the same cross-section ($10\times 10\;\mathrm{mm^{2}}$) but differ in length ($L_{1}=62\;\mathrm{mm}$ and $L_{2}=46\;\mathrm{mm}$). A representative bar is therefore defined with the averaged length $L_{\mathrm{avg}}=(L_{1}+L_{2})/2=54\;\mathrm{mm}$, giving $V_{\mathrm{sb}}=10\times 10\times 54=5400\;\mathrm{mm^{3}}$ and $A_{\mathrm{sb}}=2\,(10\times 10)+4\,(10\times 54)=2360\;\mathrm{mm^{2}}$. The single-bar sphericity is
$\psi_{\mathrm{sb}}=\frac{\pi^{1/3}\,\bigl(6\,V_{\mathrm{sb}}\bigr)^{2/3}}{A_{\mathrm{sb}}}\,$, (18)
where $V_{\mathrm{sb}}$ and $A_{\mathrm{sb}}$ are the volume and surface area of the representative square bar, respectively. For the representative bar described above, $\psi_{\mathrm{sb}}=0.63$. The Sauter-mean diameter for a single bar is
$d_{\mathrm{sd}}^{\mathrm{sb}}=\frac{6\,V_{\mathrm{sb}}}{A_{\mathrm{sb}}}\,.$ (19)
The corresponding single-bar equivalent diameter reads
$d_{\mathrm{sb}}=\psi_{\mathrm{sb}}\,d_{\mathrm{sd}}^{\mathrm{sb}}\,.$ (20)
Note that $d_{\mathrm{sb}}$ is constant for all $\alpha$, since both $\psi_{\mathrm{sb}}$ and $d_{\mathrm{sd}}^{\mathrm{sb}}$ are determined by the fixed bar geometry, yielding $d_{\mathrm{sb}}=8.65\;\mathrm{mm}$. The Blake–Kozeny permeability estimate (Eq. 13) can be evaluated with each characteristic diameter:
$K_{\mathrm{eq}}=\frac{d_{\mathrm{eq}}^{2}}{150}\,\frac{\phi^{3}}{(1-\phi)^{2}}\,,\qquad K_{\mathrm{sb}}=\frac{d_{\mathrm{sb}}^{2}}{150}\,\frac{\phi^{3}}{(1-\phi)^{2}}\,.$ (21)
For both diameter definitions, the Ergun-type Reynolds number (Eq. 12) is formed with the respective equivalent diameter:
$\mathrm{Re}_{\mathrm{Erg}}^{d_{\mathrm{eq}}}=\frac{\langle w\rangle\,d_{\mathrm{eq}}}{\nu}\,\frac{\phi}{1-\phi}\,,\qquad\mathrm{Re}_{\mathrm{Erg}}^{d_{\mathrm{sb}}}=\frac{\langle w\rangle\,d_{\mathrm{sb}}}{\nu}\,\frac{\phi}{1-\phi}\,.$ (22)
3.2 Tortuosity-based permeability model
The Blake–Kozeny and Kozeny–Carman models can be both expressed in a general sense as
$K=\frac{\phi^{3}}{c_{\mathrm{KC}}\,\tau^{2}\,a_{v}^{2}}\,$, (23)
where $c_{\mathrm{KC}}$ is the Kozeny constant (for a cylindrical capillary $c_{\mathrm{KC}}=2$) and $\tau$ is the Hydraulic tortuosity which quantifies the elongation of fluid paths relative to the straight-line distance through the medium. For spheres, $\tau$ is typically fixed at $\tau=\sqrt{2}$ (clennell1997; carman1937) leading to the Blake–Kozeny model.
Naturally, there is no certainty that a model formulated with a constant value of $\tau$ will be valid across many geometries and estimating permeability by the use of Eq. 23 may yield improved results. Different definitions of tortuosity have been proposed in the literature (bear1972; clennell1997; matyka2008; duda2011). Following duda2011, the upper bound for the hydraulic tortuosity can be computed directly from the velocity field as the ratio of the volume-averaged velocity magnitude to the volume-averaged axial velocity component in the void space:
$\tau_{\mathrm{eff}}=\frac{\langle|\bm{u}|\rangle}{\langle w\rangle}.$ (24)
At the larger Reynolds numbers, the upper bound for the hydraulic tortuosity depends on $\mathrm{Re}_{\mathrm{p}}$ through the rearrangement of the flow field by inertial effects and is denoted $\tau_{\mathrm{eff}}$. In the Darcy limit, it reduces to
$\tau=\lim_{\mathrm{Re}_{\mathrm{p}}\to 0}\tau_{\mathrm{eff}}\,$, (25)
which is a purely geometric property of the void space.
4 Numerical model
All simulations were performed using OpenFOAM-12 (greenshields2024), an open-source finite volume-based Computational Fluid Dynamics (CFD) software. Both steady-state and transient solvers were used, employing the SIMPLE and PISO pressure-velocity coupling methods, respectively (moukalled2016). To determine the appropriate solution method for each geometry and flow regime, both algorithms were tested to compute the flow at $\mathrm{Re}_{\mathrm{p}}=50$ in each geometry. The results showed no significant differences in the microscopic and macroscopic flow features, such as the pressure gradient and the intrinsic average of velocity indicating that the steady-state solver is an appropriate choice for this Reynolds number. Expecting the onset of unsteady flow for higher Reynolds numbers, all simulations with $\mathrm{Re}_{\mathrm{p}}>50$ were performed using the unsteady solver, otherwise a steady-state solver was used.
Table 1: The information regarding each of the considered geometries and computational meshes. The table contains the number of cells per layer ($N_{\mathrm{C}}$) and the number of layers ($N_{\mathrm{L}}$) for each of the considered rotation angles $\alpha$.
$\alpha$ $0^{\circ}$ $5^{\circ}$ $10^{\circ}$ $15^{\circ}$ $20^{\circ}$
$N_{\mathrm{L}}$ 1 36 18 12 9
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1 1.46 1.33 1.22 1.46
$\alpha$ $25^{\circ}$ $30^{\circ}$ $35^{\circ}$ $40^{\circ}$ $45^{\circ}$
$N_{\mathrm{L}}$ 36 6 36 9 4
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.21 1.27 1.30 1.27 1.24
$\alpha$ $0^{\circ} 55$$5^{\circ}$ $60^{\circ}$ $65^{\circ}$ $70^{\circ}$
$N_{\mathrm{L}}$ 18 36 3 36 18
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.22 1.24 1.26 1.28 1.27
$\alpha$ $75^{\circ}$ $80^{\circ}$ $85^{\circ}$ $90^{\circ}$
$N_{\mathrm{L}}$ 12 9 36 2
$N_{\mathrm{C}}$ ($\times 10^{6}$) 1.27 1.26 1.25 1.26
The residual convergence tolerance was set to $1\times 10^{-13}$ for pressure and velocity equations in steady-state simulations to ensure proper convergence of velocity gradients and accurate computation pressure drop over the packed bed geometry. The maximum number of correctors for mesh non-orthogonality was set to 3.
The unsteady simulations employed a less strict convergence tolerance equal to $1\times 10^{-8}$. The Courant number was kept below $0.8$. The number of PISO corrector steps was set to 4, with one additional loop for mesh non-orthogonality.
To force the flow, the pressure gradient momentum source was adopted, with the value calibrated during runtime, ensuring desired the bulk velocity and Reynolds number. Second-order accurate spatial discretization schemes using linear interpolation were used for both convective and diffusive terms. For time integration a second-order implicit backward scheme was adopted.
The native geometric–algebraic multigrid (GAMG) solver was predominantly used for the pressure equation in both steady and unsteady simulations. However, for rotation angles of $5^{\circ}$, $10^{\circ}$, $15^{\circ}$ and $20^{\circ}$, at high $\mathrm{Re}_{\mathrm{p}}$, the preconditioned conjugate gradient (PCG) method was used instead due to convergence difficulties. The PBiCGStab method with the DILU preconditioner was applied to the velocity equation in all cases.
Each module was meshed separately, resulting in the grids for each geometry consisting of a number of separate regions, which were connected through a non-conformal coupling (NCC) mesh interface (greenshields2024), which creates new set of faces at the intersections of overlapping faces on the coupled patches. These intersection faces become part of the finite volume mesh on both sides of the interface, ensuring that the mesh is topologically closed. As a consequence, conservation of mass and momentum is maintained by the standard finite volume discretisation, without any additional correction. The coupling is fully conservative and formally second-order accurate.
As with standard unstructured finite volume discretisation such boundary condition can still result in a decrease of the overall accuracy or convergence order, resulting from potentially high non-orthogonality and skewness of coupled faces, requiring detailed monitoring of the NCC regions. For some of the considered geometries, we have observed that the use of NCC conditioners led to the presence of pressure oscillations near the coupled boundaries. As their presence did not impede convergence and subsequent refinements near the coupled regions did not contribute to a significant change in averaged results (e.g., relative change of friction factor below 2%), we have deemed the accuracy offered by NCC acceptable.
Side walls of the module were treated as no-slip boundaries with standard zero Neumann boundary conditions for pressure. The top and bottom boundaries of each geometry were also treated using NCC interface, imposing periodic boundary conditions to replicate an infinitely repeating porous structure.
An unstructured mesh was employed for all cases, generated using Gmsh (geuzaine2009), with prism cells near the corners of smaller void spaces.
Figure 4: (a) Schematic representation of the simulation domain for the case with a module angle of $50\text{\,}\mathrm{\SIUnitSymbolDegree}$. Part of the domain is clipped for clarity. The first zoomed-in view illustrates the orientation of the mesh, while the second highlights the mesh structure between adjacent layers. (b) Detailed view of the refined mesh at the same angle
Sensitivity of the results to the chosen resolution was evaluated using three different meshes without refinement (see Appendix A for details). For the steady-state simulations, a medium-resolution mesh comprising approximately 1 million cells per module was selected. Depending on the rotation angle, this configuration produced a maximum non-orthogonality of $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ (with an average of $9.55\text{\,}\mathrm{\SIUnitSymbolDegree}$) and a maximum skewness of 0.84. As an example, a mesh with a rotation angle of $50\text{\,}\mathrm{\SIUnitSymbolDegree}$ is shown in Fig. 4.
To accurately capture the unsteady flow, for transient simulations, the meshes were refined near the overlapping parts of the mesh interfaces between the regions. The resulting increase in cell count was geometry dependent: for the $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ case, this amounted to approximately 20%, whereas for the $5\text{\,}\mathrm{\SIUnitSymbolDegree}$ case, the increase was around 50% due to the larger contact surface. After refinement, the maximum non-orthogonality increased to about $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ (average of $12.28\text{\,}\mathrm{\SIUnitSymbolDegree}$), and the maximum skewness 1.34. A representative refined mesh for $50^{\circ}$ angle case is shown in Fig. 4. Table 1 presents the number of cells for each rotated angle after the refinement.
Table 2: Normalized simulations times for different Reynolds numbers. Here, $T_{B}=B/\langle w\rangle$ denotes the characteristic time scale, $T_{\mathrm{i}}$ represents the start-up time, and $T_{\mathrm{avg}}$ the average time.
$\mathrm{Re}_{\mathrm{p}}$ $T_{\mathrm{i}}/T_{B}$ $T_{\mathrm{avg}}/T_{B}$
100 20 10
150 30 15
200 40 20
For transient simulations, the averaging time was defined such that the first-order quantities for each Reynolds number reached a statistically steady-state at randomly selected probe locations within the simulation domain. The times are listed in Table 2.
4.1 Periodic averaging & data reduction
Figure 5: Computational mesh of the $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ simulation domain, representing periodically repeated module used for averaging
For each rotation angle $\alpha$, the numerical domain consists of identical modules, repeated several times as described in Section 2. The flow field in the void spaces of each module should be identical when viewed in a reference frame rotated with the module. This allows to further boost the statistical convergence of time-averaged data by performing a periodic average, or averaging the flow field over each module in the considered geometry.
This is achieved by first identifying the corresponding cells and faces, as illustrated in Fig. 5. Scalar fields can then be directly averaged by summing the affiliated field values at these cell/faces and dividing by the total number of modules. In the case of vector and tensor fields, before averaging, the field has to be first rotated appropriately for each module. The average values can be distributed to the cells/faces corresponding to each other (applying the reverse transformation for the vector and tensor fields). Both the velocity and pressure fields are averaged this way in this study.
Finally, since the periodically averaged fields are identical across layers (without considering the rotation), only one representative module needs to be considered for analysis and data archiving. The boundary values of the representative module were calculated by interpolating between adjacent cells. As an example of the storage efficiency achieved, for $\alpha=$5\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=200$, the simulation comprising 36 layers originally required about 15 GB of storage, which was reduced to approximately 135 MB after averaging. We would like to remark that such method cannot be used for archiving of snapshots of unsteady velocity or pressure, and we employ it only for time-averaged data or out of steady-state simulations. Unless otherwise stated, the results presented in the following sections are based on periodically averaged mean values.
5 Results and discussion
5.1 Validation
Figure 6: Comparison of normalised streamwise velocity fields $\overline{w}/\langle w\rangle$ obtained from PIV measurements and CFD simulations at $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ for $\mathrm{Re_{p}}=100$ and $200$. (a) Position P1, (b) position P3.
The numerical setup was validated by comparing results for $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=100,200$ against experiments conducted by christin2025. We compared the mean flow fields at two planes located between the bars (P1 and P3), as shown in Fig. 1. The experimental configuration consists of 18 modules and an outlet section, with measurements available at layers 13th–17th and above the bed. Because the velocity at the 13th layer is least affected by outlet effects, it was selected for comparison with the fully developed, periodic flow of the simulation.
The flow field at P1 and P3 is compared in Fig. 6, plotted in a $(\xi,z)$ coordinate system (horizontal and vertical coordinates, respectively). The simulation captures the experimental flow features with satisfactory accuracy, particularly at the midspan region. Both datasets indicate that the flow field behaviour is mainly determined by the geometry of the void space, especially by the locations of the inlets and outlets. The accelerated fluid near these narrow connections between the modules generates several recirculation regions attached to the bottom and top walls.
While the midspan velocity distributions show strong agreement in both flow direction and normalized velocity, deviations are observed at the far left and right sides of the domain. These discrepancies are attributed to the experimental setup: the RC2 geometry required modification to allow for optical access. For a comprehensive comparative analysis of the numerical and experimental results, the reader is referred to sadowski2025.
Figure 7: Normalised spanwise velocity profiles ($\overline{w}/\langle w\rangle$) from PIV measurements and CFD simulations at $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Profiles are shown at three vertical positions (z/B=12.25, 12.50, 12.75z/B=12.25,\,12.50,\,12.75) for $\mathrm{Re_{p}}=100$ and $200$.
Figure 7 presents the spanwise velocity profiles normalized by the intrinsic velocity at three $z$-positions of the 13th layer for $\mathrm{Re}_{\mathrm{p}}=100$ and $200$.
At $\mathrm{Re}_{\mathrm{p}}=100$, the simulations reproduce the characteristic flow features with high fidelity, capturing both the location and magnitude of the peak velocities. This quantitative agreement is confirmed by the root mean square deviation (RMSD) analysis, which yielded its lowest value (0.333) at this Reynolds number at location P3. However, three notable deviations are observed. First, minor discrepancies appear near the lateral boundaries, particularly at z=12.75mmz=12.75\ $\mathrm${mm}. Second, a boundary-layer jet is clearly resolved at P3 in the CFD data but is essentially absent in the PIV measurements; both are likely due to the geometric differences in the experimental setup discussed previously. Third, at the far left and right sides, P3 exhibits a jet-like flow while P1 does not (a trend also observed at $\mathrm{Re}_{\mathrm{p}}=200$). This is attributable to the location of P3, which is closer to the central axis, at this orientation.
At $\mathrm{Re}_{\mathrm{p}}=200$, the flow is unsteady, increased inertial effects lead to significantly stronger vortices and the emergence of distinct local velocity fluctuations. Reflecting this increased physical complexity, the RMSD at P1 is 1.009. Although this represents the maximum deviation across the simulated cases, it remains within acceptable limits considering the transient nature of the flow, and the overall agreement between the numerical and experimental results remains robust.
5.2 Overview of the flow fields
Figure 8: Normalized mean velocity magnitude without periodic averaging at $\mathrm{Re}_{\mathrm{p}}=200$ for (a) $20\text{\,}\mathrm{\SIUnitSymbolDegree}$ and (b) $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ geometries, visualized on the plane at position P3 (see Fig. 1). The flow direction is along the +z+z axis.
We provide an overview of the flow field under different geometric configurations and flow conditions. Figure 8 presents the mean velocity field for two different angles ($\alpha=$20\text{\,}\mathrm{\SIUnitSymbolDegree}$,$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) at the highest Reynolds number ($\mathrm{Re}_{\mathrm{p}}=200$), on the slices through the void spaces in each module located at the position P3 (see Fig. 1). The strong similarities observed between successive layers highlight the periodic nature of the flow, thereby justifying the use of periodically averaged quantities in the following sections.
[…]
Figure 9: Normalized mean velocity magnitude without periodic averaging in a packed bed with a rotation angle of $60\text{\,}\mathrm{\SIUnitSymbolDegree}$. Slices are taken from the middle of each module: (a) at $\mathrm{Re}_{\mathrm{p}}=1$, (b) at $\mathrm{Re}_{\mathrm{p}}=200$.
After examining the influence of the rotation angle on the flow structure, the effect of the Reynolds number on the flow regime was considered. Figure 9 shows the normalized mean velocity magnitude for a geometry with $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at two different Reynolds numbers. The slices shown in Fig. 9 are extracted from the mid-planes of three successive modules, located at $z=5$, $15$, and $25\text{\,}\mathrm{m}\mathrm{m}$, respectively.
At $\mathrm{Re}_{\mathrm{p}}=1$ (Fig. 9), the flow is fully laminar and stationary with the pressure drop primarily governed by Darcy’s law. The fluid passes smoothly through the void spaces between layers, where the normalized velocity magnitude ranges from 0 to 2.5. The streamlines follow smooth, slightly curved trajectories directed upward through the bed. The velocity distribution within the void space exhibits a low-speed, upward-directed flow that accelerates toward the center.
At $\mathrm{Re}_{\mathrm{p}}=200$ (Fig. 9), the flow is in the inertia-dominated regime. In contrast to $\mathrm{Re}_{\mathrm{p}}=1$ (Fig. 9), the streamlines here exhibit significant tortuosity. This complexity arises because the increased inertial forces prevent the fluid from following the curvature of the solid bars, leading to the formation of low-velocity wake regions and distinct recirculation zones within the void spaces. The flow develops strong transverse velocity components, which drive significant deviations from the primary streamwise direction.
5.3 Friction factor
Figure 10: Friction factor ($f_{\mathrm{p}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different rotation angles
Figure 11: Friction factor ($f_{\mathrm{eq}}$) as a function of $\mathrm{Re}_{\mathrm{eq}}$ for different rotation angles. The dotted black line shows the Ergun correlation (Eq. 11). Inset: same data computed with the single-bar diameter $d_{\mathrm{sb}}$.
Figure 10 shows the relation of the friction factor as a function of Reynolds number for various angles. Each line represents the friction factor calculated from Eq. 6. The color gradient indicates the rotation angle of each geometry from $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$.
In the viscous regime, all curves follow a hyperbolic trend with respect to $\mathrm{Re}_{\mathrm{p}}$, which appears as a straight line in the log–log scale. As the angle increases from $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $25\text{\,}\mathrm{\SIUnitSymbolDegree}$, the friction factor increases and reaches its maximum at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Beyond this angle, the friction factor decreases gradually, and further changes in rotation angle no longer introduce a large variation in $f_{\mathrm{p}}$.
As the Reynolds number increases, deviations from the linear trend become apparent, signaling the onset of inertial effects and the transition toward the non-linear (Forchheimer) regime, for larger angles. However, the configuration at $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ exhibits a much slower transition to the inertial regime (see Section 5.6) with $f_{\mathrm{p}}$ resembling a straight line. Additionally, the angle yielding the maximum friction factor shifts with the Reynolds number. For instance, at $\mathrm{Re}_{\mathrm{p}}=200$, the maximum friction factor occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, in contrast to the peak at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ observed in the viscous regime.
These values of the drag can be traced to the properties of each individual geometry. The specific alignment of the square bars at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ creates a severe constriction of the flow path. These narrowing void spaces increase the local shear stress, resulting in the maximum friction factor observed in the viscous regime. At $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the flow shows behaviour similar to contraction–expansion sequences in ducts. As the Reynolds number increases, the acceleration into void spaces creates local jets, leading to flow separation and recirculation downstream of these constrictions. Consequently, form drag becomes the dominant loss mechanism, resulting in the highest observed friction factor. In the case of $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the perfect orthogonality between modules allows the flow to develop with minimal obstruction, resulting in reduced flow resistance. Despite this low friction factor, the geometric parameters of the $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration ($D/d_{\mathrm{eq}}=3.23$, $a_{v}=0.193~\mathrm{mm^{-1}}$) place it within the lattice-like group rather than the channel-like regime observed at $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
The accuracy of the Ergun depends strongly on the choice of characteristic length scale. Figure 11 presents the friction factor computed with the module-equivalent diameter $d_{\mathrm{eq}}$. In the viscous regime, the curves for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) collapse onto the Ergun correlation. This is attributed to the fact that $d_{\mathrm{eq}}$ incorporates the angle-dependent wetted surface area through $a_{v}$ (Eqs. 17 and 8), unlike $d_{\mathrm{sb}}$, which remains constant for all rotation angles. With increasing Reynolds number, the curves progressively diverge: at intermediate angles (notably $\alpha\approx$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) the friction factor exceeds the Ergun prediction, whereas the channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ remain below it.
When the constant single-bar diameter $d_{\mathrm{sb}}$ is employed instead (inset of Fig. 11), the Ergun correlation overestimates the friction factor systematically across all angles and Reynolds numbers. Furthermore, the spread among the individual curves is considerably larger, since $d_{\mathrm{sb}}$ does not vary with $\alpha$ and is therefore unable to capture the differences in wetted surface area between configurations. These observations indicate that $d_{\mathrm{eq}}$ is the more suitable length scale for correlating the friction factor in the present geometry.
5.4 Determining permeability and forchheimer coefficient
The determination of the $\mathrm{Da}$ and $C_{\mathrm{F}}$ numbers was carried out in two steps using a least-squares fitting optimization procedure of Eq. 7 to the gathered friction factor data. First, the Darcy number was obtained for $\mathrm{Re}\leq 1$ by neglecting inertial effects (i.e. assuming $C_{\mathrm{F}}=0$). In the second step, the $C_{\mathrm{F}}$ was evaluated by fitting the Eq. 7 over the entire Reynolds number range while keeping the previously determined Darcy number fixed.
Figure 12 presents a comparison between the computed friction factor ($f_{\mathrm{p}}$), obtained from Eq. 6, and the fitted non-dimensional friction factor ($f_{\mathrm{fitted}}$). The latter was evaluated using Eq. 7, incorporating the calculated Darcy number and Forchheimer coefficient. As shown in Fig. 12, the fitted curves closely follow the computed data over the selected range of Reynolds numbers. The agreement is strong in the viscous regime, while slight deviations appear at higher Reynolds numbers, suggesting that additional simulations at higher $\mathrm{Re}_{\mathrm{p}}$ may be required to better capture the inertial effects.
The computed Darcy number, plotted in Fig. 13, shows a strong dependence on the rotation angle, demonstrating two distinct behaviours depending on the geometry type. For $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the gaps in consecutive modules are aligned in a channel-like form, which promotes a less perturbed flow and results in high permeability. As the angle increases, the connections between consecutive modules become more obstructed, leading to a sharp decrease in permeability with $\alpha$, observable also as a previously described trend in the friction factor in the viscous regime.
[…]
The configurations at $\alpha=$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $20\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ are classified as lattice-like based on their bed-to-particle diameter ratio ($D/d_{\mathrm{eq}}>3$). This classification is confirmed by their Darcy numbers ($\mathrm{Da}=0.0020$, $0.0019$ and $0.0018$, respectively), which fall within the range of the lattice-like group ($\mathrm{Da}=0.0017\pm 0.0004$). For the $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration in particular, $D/d_{\mathrm{eq}}=3.23$ and $a_{v}=0.193~\mathrm{mm^{-1}}$ are consistent with the lattice-like geometries, in contrast to the channel-like configurations at $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, where $D/d_{\mathrm{eq}}<3$ and $a_{v}<0.16~\mathrm{mm^{-1}}$. This confirms that the bed-to-particle diameter ratio ($D/d_{\mathrm{eq}}\approx 3$ as boundary) and the hydraulic tortuosity ($\tau\approx 1.1$, see Fig. 14) provide a reliable basis for the classification.
Figure 12: Comparison between the computed friction factor ($f_{\mathrm{p}}$) and the fitted friction factor ($f_{\mathrm{fitted}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different angles. Scaling factors are applied to separate the curves vertically for clarity.
Figure 13: Darcy–Forchheimer model parameters as a function of the rotation angle $\alpha$. (a) Darcy number $\mathrm{Da}=K/B^{2}$, the tortuosity-based model ($K_{\tau}$) and Blake–Kozeny correlations using $d_{\mathrm{eq}}$ and $d_{\mathrm{sb}}$. The dashed line and shaded band indicate the DNS average ($0.0017$) and $\pm 2\sigma$ ($0.0004$), respectively. (b) Forchheimer coefficient $C_{\mathrm{F}}$ obtained from DNS fitting compared with values computed using $d_{\mathrm{sb}}$ and $d_{\mathrm{eq}}$.
The Forchheimer coefficient $C_{\mathrm{F}}$, shown in Fig. 13, exhibits a trend that parallels the Darcy number at small rotation angles: channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) yield the lowest values of $C_{\mathrm{F}}$. This is not a direct consequence of the high permeability at these angles, but rather reflects a common geometric origin: aligned void spaces suppress flow separation and recirculation, thereby reducing both viscous resistance (high $\mathrm{Da}$) and inertial losses (low $C_{\mathrm{F}}$).
For $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the low $C_{\mathrm{F}}$ is consistent with the delayed transition to the inertial regime observed for this configuration (see Section 5.6), which suggests that the Forchheimer term is not fully developed within the present Reynolds number range.
At intermediate angles ($30\text{\,}\mathrm{\SIUnitSymbolDegree}$$70\text{\,}\mathrm{\SIUnitSymbolDegree}$), $C_{\mathrm{F}}$ exhibits considerably stronger variations with $\alpha$ than at the extremes. In this range, the void-space topology changes significantly with each increment in rotation angle, producing markedly different levels of flow acceleration, separation and recirculation.
5.5 Modelling permeability and forchheimer coefficient
Figure 14: (a) $\tau_{\mathrm{eff}}$ (Eq. 24) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different rotation angles. (b) Hydraulic tortuosity $\tau$ (Eq. 25) as a function of the rotation angle $\alpha$.
Figure 14(a) shows the upper bound for the hydraulic tortuosity $\tau_{\mathrm{eff}}$, computed from Eq. 24, as a function of the Reynolds number for all rotation angles. In the viscous regime, $\tau_{\mathrm{eff}}$ remains approximately constant for each angle, consistent with the Darcy flow assumption where the flow topology is independent of the flow rate. As the Reynolds number increases, inertial effects alter the flow paths and $\tau_{\mathrm{eff}}$ increases for most angles. The channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) exhibit the lowest values ($\tau_{\mathrm{eff}}\approx 1$), reflecting nearly straight flow paths. The $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration also shows a low tortuosity ($\tau_{\mathrm{eff}}\approx 1.09$), placing it at the lower bound of the lattice-like group.
Figure 14(b) presents the hydraulic tortuosity $\tau$ (Eq. 25) as a function of the rotation angle. For $\alpha\leq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\tau$ is close to unity, indicating that the aligned bar arrangement allows the flow to pass with minimal deviation. A sharp increase occurs between $\alpha=$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ and $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, coinciding with the transition from channel-like to lattice-like geometries. For $\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\tau$ oscillates between approximately $1.1$ and $1.35$, with a peak near $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. This is consistent with the friction factor results (Fig. 10), where the maximum drag in the viscous regime was also observed at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
Figure 13 compares the DNS permeability with three model predictions. The Blake–Kozeny estimate based on the module-equivalent diameter, $K_{\mathrm{eq}}$ (Eq. 21), provides the closest agreement with the DNS data for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), with a mean absolute percentage error (MAPE) of $12.3\%$. Because $d_{\mathrm{eq}}$ incorporates the angle-dependent wetted surface area through $a_{v}$ (Eqs. 8 and 10), it captures the dominant geometric effect governing permeability variations in these configurations.
The tortuosity-based estimate $K_{\tau}$ (Eq. 23), which combines the hydraulic tortuosity $\tau$ with the specific surface area $a_{v}$ and a fitted Kozeny constant $c_{\mathrm{KC}}=3.1$, reproduces the angle-dependent trend with a MAPE of $18.9\%$. Both $K_{\mathrm{eq}}$ and $K_{\tau}$ provide good estimates of the permeability for the lattice-like geometries, with MAPEs below $20\%$. In contrast, $K_{\mathrm{sb}}$, based on the constant single-bar diameter, is nearly independent of $\alpha$ and yields a MAPE of $80.2\%$, confirming that it cannot reproduce the angle-dependent permeability.
For the channel-like geometries ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), all three models underestimate the permeability. These correlations are derived for multiply connected pore structures and do not account for the aligned, low-resistance void spaces present at small rotation angles.
The Ergun-based Forchheimer coefficients, $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{sb}})$ and $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{eq}})$, computed from Eq. 14 using $K_{\mathrm{sb}}$ and $K_{\mathrm{eq}}$ respectively, are compared with the DNS-fitted values in Fig. 13. Because Eq. 14 relates $C_{\mathrm{F}}$ directly to $\sqrt{K}$, the quality of the permeability estimate propagates into the predicted Forchheimer coefficient.
The single-bar estimate $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{sb}})$ yields a constant value of approximately $0.7$, which underestimates the DNS-fitted $C_{\mathrm{F}}$ for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$). Since $d_{\mathrm{sb}}$ does not vary with the rotation angle, the resulting $K_{\mathrm{sb}}$ cannot capture the angle-dependent permeability, and consequently the predicted Forchheimer coefficient inherits this limitation. The module-equivalent estimate $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{eq}})$ varies with $\alpha$ through the angle-dependent specific surface area and reaches values closer to the DNS data at some intermediate rotation angles. However, it significantly overestimates $C_{\mathrm{F}}$ at small rotation angles ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), where the channel-like flow structure departs from the packed-bed assumptions underlying the Ergun correlation. At $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, both Ergun estimates also overestimate $C_{\mathrm{F}}$, consistent with the low DNS-fitted value discussed above, which is attributed to the delayed transition to the inertial regime at this angle. Moreover, the DNS data are limited to $\mathrm{Re}_{\mathrm{p}}\leq 200$, which may not be sufficient to fully establish the Forchheimer coefficient for all geometries. The development of geometry-aware models for the Forchheimer coefficient, supported by simulations at higher Reynolds numbers, remains a subject for future investigation.
5.6 Transition to inertial regime
Figure 15: Relative deviation of the friction factor from the Darcy-regime value. The circular markers with black outlines show the critical Reynolds number for each angle.
According to the classification proposed by Dybbs1984ANL, this transition typically occurs near $\mathrm{Re}_{\mathrm{p}}\approx 1$ for spherical particles in packed beds. The present geometry consists of square bars arranged at a constant rotation angle, leading to void space connectivity characteristics that differ significantly from those of a sphere-packed medium.
Hlushkou2006 suggested that if the inertial contribution exceeds $5\%$, the flow should no longer be considered in the Darcy regime. The same criterion is adopted in the present study. To identify the Reynolds number at which inertial effects become significant, the difference between the total friction factor and the linear friction factor ($f_{\mathrm{lin}}=1/(\phi\,\mathrm{Re}_{\mathrm{p}}\,\mathrm{Da})$) is computed and then divided by the total friction factor. However, other criteria have also been reported in the literature, such as introduced by Zeng2006; bagci2014. In future work, different transition criteria may be used to provide a more comprehensive assessment of the transition to inertia-dominated flow.
We illustrate the transition from viscosity to inertia-dominated regime in Fig. 15. The upper figure shows the relative friction factor for each angle. As expected, the deviation remains zero for $\mathrm{Re}_{\mathrm{p}}\leq 1$. With increasing $\mathrm{Re}_{\mathrm{p}}$, the inertial contribution represented by the Forchheimer term grows with the Reynolds number. Consequently, the pressure drop no longer follows the linear Darcy regime, and the deviation from Darcy’s law increases rapidly.
[…]
At $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the transition to the inertial regime occurs at approximately $\mathrm{Re}_{\mathrm{p}}\approx 5$, which is lower than for the other geometries. Potential factors contributing to this behaviour include the sensitivity of the Darcy number evaluation for this orientation and the unique connectivity of the void space, which may promote early localized recirculation. A conclusive explanation would require a detailed flow field analysis which currently exceeds the scope of this study.
6 Conclusion
In this study, numerical simulations were performed using OpenFOAM-12 to investigate viscosity- and inertia-dominated flow through periodically arranged square-bar structures over a range of $\mathrm{Re}_{\mathrm{p}}=0.1\text{--}200$ with different angles ($0\text{\,}\mathrm{\SIUnitSymbolDegree}$$90\text{\,}\mathrm{\SIUnitSymbolDegree}$) at fixed porosity $\phi=0.332$. A total of 266 simulations were performed using a predominantly hexahedral mesh, and selected cases were successfully validated against available PIV measurements obtained from the experimental configuration of christin2025. Some important conclusions can be drawn from the obtained results:
  1. (1)
    The rotation angle has a strong influence on the macroscopic flow characteristics. For $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structure forms channel-like geometries, resulting in high permeability. As the angle increases beyond $10^{\circ}$, the connectivity between successive layers becomes more complex, leading to the development of recirculation zones and increased frictional losses, resulting in heterogeneous velocity fields. We refer to these configurations ($\alpha\geq 15^{\circ}$, including $90^{\circ}$) as lattice-like geometries. The $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration is also classified as lattice-like, as confirmed by its Darcy number, bed-to-particle diameter ratio and tortuosity.
  2. (2)
    Using a 5% inertial contribution threshold, the transition to the inertial regime occurs at $\mathrm{Re}_{\mathrm{p}}\approx 7.5$ for lattice-like geometries, with the earliest onset at $\alpha=$55\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
  3. (3)
    The drag characteristics are strongly impacted by the geometric configuration, and the angle yielding the highest friction factor shifts with the Reynolds number. In the viscous regime, the severe constriction of the flow path at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ results in the highest friction factor. In the inertial regime, the contraction–expansion sequences at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ lead to flow separation and recirculation, resulting in the highest friction factor.
  4. (4)
    The classification into channel-like and lattice-like regimes can be supported quantitatively by the bed-to-particle diameter ratio and the hydraulic tortuosity. Geometries with $D/d_{\mathrm{eq}}<3$ and $\tau<1.1$ behave as channel-like, whereas those with $D/d_{\mathrm{eq}}>3$ and $\tau>1.1$ fall into the lattice-like group.
  5. (5)
    The module-equivalent diameter $d_{\mathrm{eq}}$, which accounts for the angle-dependent wetted surface area, collapses the friction factor data onto the Ergun correlation more effectively than the constant single-bar diameter $d_{\mathrm{sb}}$, allowing to partially account for the geometry variation. Among the permeability models, the Blake–Kozeny estimate based on $d_{\mathrm{eq}}$ provides the best agreement with the DNS data for the lattice-like geometries (MAPE $12.3\%$). The tortuosity-based estimate $K_{\tau}$ (Kozeny constant $c_{\mathrm{KC}}=3.1$) reproduces the angle-dependent trend but with larger deviations (MAPE $18.9\%$), while the single-bar estimate yields substantially larger errors (MAPE $80.2\%$). All three models underestimate the permeability of the channel-like configurations ($\alpha\leq 10^{\circ}$), as they are formulated for multiply connected pore structures.
  6. (6)
    The Forchheimer coefficient $C_{\mathrm{F}}$ varies strongly with rotation angle, peaking at $\alpha=60^{\circ}$ and remaining low for the channel-like and $90^{\circ}$ configurations. At intermediate angles ($30^{\circ}$$70^{\circ}$), $C_{\mathrm{F}}$ fluctuates most strongly, reflecting the sensitivity of inertial losses to the void-space topology in this range.
The good agreement between simulations and experiments confirms that the present approach provides a reliable framework for analysing flow through such periodic structures. Future work should extend the analysis to higher Reynolds numbers to examine the onset of unsteady and turbulent flow, and include additional configurations to improve the fitting of geometry-aware models for both permeability and the Forchheimer coefficient.
Acknowledgments
The authors gratefully acknowledge the financial support of the Deutsche Forschungsgemeinschaft (DFG) through No. SFB/ TRR287, Project No. 422037413. The authors wish to thank Christin Velten, Kerstin Hülz, and Katharina Zähringer for sharing the dataset of the experimental measurements of flow in the packed bed.
Appendix A Effect of mesh resolution on the numerical results
Figure A.1: Normalized spanwise-averaged velocity profiles at P3 (see Fig. 1) for angles $30\text{\,}\mathrm{\SIUnitSymbolDegree}$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. Results are shown for three mesh resolutions (200K, 1M, and 4M cells)
Figure A.2: Friction factor for three mesh sizes (200K, 1M, and 4M cells) at angles of $30^{\circ}$, $60^{\circ}$, and $90^{\circ}$.
Table A.1: Friction factors used in for each angle and mesh. The table contains the number of cells per layer ($N_{\mathrm{C}}$)
$\alpha$ $N_{\mathrm{C}}$ $2\text{\times}{10}^{5}$ $1\text{\times}{10}^{6}$ $4\text{\times}{10}^{6}$
$30\text{\,}\mathrm{\SIUnitSymbolDegree}$ 67.70 70.27 72.87
$60\text{\,}\mathrm{\SIUnitSymbolDegree}$ 85.33 87.84 91.01
$90\text{\,}\mathrm{\SIUnitSymbolDegree}$ 16.36 14.53 14.47
A mesh sensitivity analysis to the results was performed using three unrefined meshes containing approximately $2\text{\times}{10}^{5}$, $1\text{\times}{10}^{6}$, and $4\text{\times}{10}^{6}$ cells per module at $\mathrm{Re}_{\mathrm{p}}=200$. No local refinement was applied in the flow region. Three modules with different rotation angle, namely $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $60\text{\,}\mathrm{\SIUnitSymbolDegree}$, and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, were considered. The objective of this analysis was to assess the influence of the overall mesh resolution on the solution accuracy and to determine an optimal mesh for the subsequent simulations.
The analysis was conducted for the case shown in Fig. A.1, focusing on the first module of the computational domain along the probe line located at point P3 (see Fig. 1), positioned at the mid-line in the $z$-direction. Velocity data were extracted from statistically steady-state solutions without applying periodic averaging. Figure A.1 presents the normalized spanwise-averaged velocity profiles. The coarsest mesh ($2\text{\times}{10}^{5}$ cells) shows a noticeable deviation from the finer ones, particularly in the central region of the $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ case. For $60\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the deviation becomes evident near the peak locations of the velocity profiles, where the $2\text{\times}{10}^{5}$ cell mesh results in less accurate results. As the mesh is refined, the profiles converge toward those of the finest mesh, indicating improved resolution of the flow field. For all investigated angles, the $1\text{\times}{10}^{6}$ and $4\text{\times}{10}^{6}$ meshes yield nearly identical profiles, suggesting that the flow features are well-resolved from the $1\text{\times}{10}^{6}$ cell mesh.
[…]